OpenStudy (anonymous):

a 6,000.00 principal earns 8% interest compounded annually. after 35 years what is the balance in the account

OpenStudy (luigi0210):

Do you know the formula?

OpenStudy (anonymous):

no but this is what i tryed. 6,000.00 times 35 and u get 210000 then i divided that by 8% and i got 16800

OpenStudy (anonymous):

but thats not right

OpenStudy (tkhunny):

Simple interest! That is not what the question asks.

OpenStudy (anonymous):

what do u mean

OpenStudy (tkhunny):

You seem to be saying that you do NOT know any formula for Compound Interest, as you have just attempted a Simple Interest formula to solve this problem.

OpenStudy (anonymous):

i guess so

OpenStudy (anonymous):

can you just tell me how to get the answer and ill be on my way

OpenStudy (tkhunny):

I'm never in a rush. Over a period of one year, Simple and Compound are exactly thee same. Simple: 6000*(1 + 0.08) = 6480 Compound: 6000*(1 + 0.08) = 6480 Anything else comes out differently. Two years: Simple: 6000 + 480*2 = 6000 + 960 = 6960 Compound: 6000(1.08)^2 = 6000*1.1664 = 6998.4 Ten years: Simple: 6000 + 480*10 = 6000 + 4800 = 10800.00 Compound: 6000(1.08)^10 = 6000*2.158925 = 12953.55 Can you do 35 years?

OpenStudy (anonymous):

where did you get 2.15892

OpenStudy (tkhunny):

\(1.08^{10} = 2.158925\)

OpenStudy (anonymous):

ok so 6000(1.08)^35 = 6000 * 14.78534 = 8871204000

OpenStudy (anonymous):

am i right??

OpenStudy (tkhunny):

You're missing a decimal place in there, somewhere.

OpenStudy (anonymous):

ive got another one for u??

OpenStudy (tkhunny):

Two minutes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a 4,000.00 principal earns 5% interest compounded annually. after 4 years what is the balance in the account

OpenStudy (anonymous):

i tried the same thing as we didi up there but the answer i got isnt an answer choice

OpenStudy (tkhunny):

That is the exact same problem. Change the numbers and do it! Hurry up. One more minute.

OpenStudy (anonymous):

i tried

OpenStudy (anonymous):

but the answer choice isnt on there

OpenStudy (tkhunny):

What did you get? "I tried" doesn't tell me anything.

OpenStudy (anonymous):

5283900

OpenStudy (tkhunny):

You DO have to start using decimal points properly. The answer absolutely is NOT 5283900. What interest rate did you use? What number of years did you use? 4000 (1.05^4) = ??

OpenStudy (anonymous):

i diid this 1.05^4 and got 4.2 then multipled that by 4000

OpenStudy (anonymous):

16800

OpenStudy (anonymous):

but thats not an answer choice

OpenStudy (tkhunny):

Did you try $4,679.43 My guess is that you need to get far more acquainted with your calculator and your formula.